# Privacy in Speech Technology

### v2

@@ I Introduction
Speech is a mode of communication and thus inherently contains a wealth of information (see table I). Communicating information already known by the receiver is pointless, and efficient communication will thus mainly contain information that is not widely known or which is private. In addition, speech contains also a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information. Finally, speaking is a dynamic interaction between two or more speakers. Dialogues thus also contain information about the relationship between participants such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history. From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled, summarized by and reflecting the personal judgment of the author.
@@
Biological Body characteristics Sustained No
State of health Variable No
Level of intoxication Variable Yes
Psychological Emotions Variable Partly
Intelligence Sustained No
Education, skill Sustained Yes
Gender identity Sustained No
@@
Variable Yes
Mannerisms, context Variable Partly
Language
choice & skills
Partly Partly
@@
Sustained Partly
Has met with person Variable Yes
@@
Variable Yes
Hardware
used
Sensor type & manufacturer Variable Yes
@@
Algorithmic stereotyping Recommender systems based on voice can become biased with respect to age, identity, religion, or ethnicity, in ways that are nearly impossible to monitor.
@@
monitoring
The subjective feeling of being continuously monitored may cause psychological damage. It may also stifle political expression damaging democratic societies.
@@
With respect to threats, media attention is often focused on breaches which have large economic consequences [Lynskey2019GuardianPrivacyScandal, hern2019, brewster2021]. While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways. First, a breach is a worst-case event, whereas threats that have not yet led to a breach event can already have a large impact. For example, users may choose not to use systems that threaten their privacy; known weaknesses and even a lax attitude of the service provider toward privacy can, therefore, have an effect on the adoption, retention, and sales of products and services. Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat. To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
@@
Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [levy2020intimate]. With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II). While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large and their prevalence makes them a considerable threat also economically.
@@
This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms. Analysis of threats, however, demonstrates that while attack surfaces have great variety, the threat models are similar (see section III). Protections against those threats have five distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, and improving overall system performance (see section IV). To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section V). While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level. For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VI). The range of applications and content types where privacy-preserving processing methods are needed is vast, and section VII presents a brief overview thereof. Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
@@
To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology. Recent works in the field, however, include a technical overview [nautsch2019preserving], as well as a popular-science review [backstrom2020privacy] of this area, and several doctoral theses have their respective summaries, e.g. [Nelus2022, das2021thesis, zarazaga2022thesis, pathak2012privacy, williams2022learning, portelo2015privacy]. Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [mahendran2021privacy], in statistical theory of privacy [cummings2023challenges], in social sciences [petronio2002boundaries] and psychology [Derlega1977PrivacyAS]. There is even an excellent and thorough meta-study of all areas of science that discuss privacy [van2018handbook].
@@
Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows). Each threat is marked with the number of the corresponding part of section III-D. Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
II Goals and Definitions
This paper focuses on the two primary uses of speech technology: telecommunication and human-computer interfaces. Concerning telecommunication approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [penders2004privacy, wong2009privacy, arapinis2017analysis]. The primary focus of this paper is thus human-computer interfaces where a computer processes speech signals (see fig. 1), but threats to communication applications are also considered where applicable. Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g. [mahendran2021privacy]).
@@ II Goals and Definitions
Privacy is closely related to challenges in security and it is often difficult to distinguish between them. Here, we strike a balance by considering privacy scenarios where an agent has legitimate access to some private speech data of the user but uses it for purposes contradicting with, or gains access beyond, the users’ expectations or preferences. For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig. 1). Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
@@
A private message is the content of a communication that is required to fulfill the purpose of said communication. Side-information is any and all information in a communication other than the private message. See fig. 2.
We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences. An malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a curious attacker seizes emerging opportunities.
@@
Attributes of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features. Table I lists examples of speech attributes with private information. Attribute or property inference refers to the identification of such private attributes of speakers of known identity or otherwise. Re-identification refers to identifying individuals from apparently anonymized information [ISObiometric, ISOIEC224745_2022].
Enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s). It is used in the context of trusted tasks, where the agent is authorized to access this information. In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data. Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data. [tomashenko2022voiceprivacy]
When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer and this is known as the trusted task. In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform tasks in a nefarious way.
The attack surface is the specific point in the system where private information leaves the authorized path to make its way to the attacker.
Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with another. Anonymization of a speech signal refers to voice conversion where the purpose is to conceal private information. Pseudonymization further refers to reversible anonymization, where the original signal can be recovered if the concealed attributes are known. [^1] These definitions of anonymization and pseudonymization are outcomes of a community-wide discussion about personally identifying information (PII) at the Lorentz Center workshop “Speech as PII: Ethics & Usability in vis-à-vis with Security & Privacy” in 2021. The anonymized speech signal then corresponds to a speech signal of an artificial person, who is known as a pseudospeaker.
III Threat and attack models
To better understand the decisive aspects of threats, it is useful to model threats and their components. The most important components of such models include the exposed data sources (section III-A), tactics used by, and data available to the attacker (section III-B), the flow of data on an abstract level, also known as the attack model (section III-C), as well as the specific point in the system where private information leaves the intended path, known as the attack surface (section III-D).
@@ III Threat and attack models
!! Private message Side-information
@@
Figure 2: The high-level threat model of speech interaction, where a private message is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message. It is a threat to privacy when an undesired recipient gains access to that private message or side information (marked by red arrows and exclamation marks).
III-A Data sources
@@
Speech data can be exposed in two ways (see fig. 2): First, when a private message is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party. Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message. Typically, speech information thus always contains consequential side information bundled into the private message. It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the private message. Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
@@
The difference between the two forms of information makes an important difference in approaches to mitigation. Transmitting the private message to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message. In some cases, though, we can reduce the amount of information content in the private message, and more generally, we should always secure the transmission for example through cryptography (see sections IV-D, IV-B and V-A2). Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task. However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section IV-A).
III-B Attack types
@@
Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [menezes2018handbook], where the attacker uses a known weakness on a large pool of potential victims. It is like spam email; if the attacker sends a large enough number of such emails, then randomly someone triggers the trap by mistake. In this form of attack, anyone could be the victim. Where motivated attacks typically target high-value victims such as famous persons where both protections and rewards are high, birthday attacks typically target ordinary people where the likelihood of success is larger but rewards are smaller.
@@
Attacks can further be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used. Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker. For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics such as linguistic information. The anonymized sentences are known as trial utterances. The attack scenarios were then classified as [tomashenko2022voiceprivacy]:

### v3

@@ I Introduction
Speech is a mode of communication and thus inherently contains a wealth of information (see table I). In addition to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information. Moreover, speaking is a dynamic interaction between two or more speakers. Dialogues thus also contain information about the relationship between participants, such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history. From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [1, 2, 3, 4].
@@
Biological Body characteristics Long-term No
State of health Short-term No
Level of intoxication Short-term Yes
Psychological Emotions Short-term Partly
Health Mixed Partly
Intelligence Long-term No
Education, skill Long-term Yes
Gender identity Long-term No
@@
Short-term Yes
Mannerisms, context Short-term Partly
Language choice Short-term Yes
Language skills Long-term Partly
@@
Mixed Partly
@@
Short-term Yes
Hardware &
software used
Sensor and software
type & manufacturer
Short-term Yes
@@
Algorithmic stereotyping Recommender systems based on voice can become biased with respect to age, identity, religion, or ethnicity in ways that are nearly impossible to monitor.
@@
monitoring, loss of
dignity
The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression damaging democratic societies. Users may feel they lose control of personal aspects of their lives.
@@
With respect to threats, media attention is often focused on breaches which have large economic consequences [5, 6, 7]. While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways. First, a breach is a worst-case event, whereas threats that have not yet led to a breach event can already have a large impact. For example, users may choose not to use systems that threaten their privacy; known weaknesses and even a lax attitude of the service provider toward privacy can, therefore, have an effect on the adoption, retention, and sales of products and services. Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat. To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
@@
Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [8]. With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II). While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large and their prevalence makes them a considerable threat also economically.
@@
This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms. The range of applications and content types where privacy-preserving processing methods are needed is vast, and section II presents a brief overview thereof, and section III defines the vocabulary. Analysis of threats demonstrates that while attack surfaces have great variety, the threat models are similar (see section IV). Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V). To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section VI). While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level. For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VII). Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
This paper focuses primarily on speech technology in human-computer interfaces (see fig. 1), but will also address telecommunication where appropriate. Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g. [9]). Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [10, 11, 12].
@@
To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology. Recent works in the field, however, include a technical overview [13], a survey on privacy issues with voice assistants [14], as well as a popular-science review [15] of this area, and several doctoral theses have their respective summaries, e.g. [16, 17, 18, 19, 20, 21, 22, 23]. Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [9], in statistical theory of privacy [24], in social sciences [25] and psychology [26]. There is even an excellent and thorough meta-study of all areas of science that discuss privacy [27]. Central contributions of this work include 1. structuring and collecting information, research results, and vocabulary into a unified format, 2. proposing a framework for analysis of attack vectors (see fig. 1), and 3. an introduction to the psychological aspects of privacy that is not often discussed in engineering-oriented publications (see section VII).
II Content categories and applications
Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy. We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g. [28, 37, 44, 44, 45, 46, 47, 48, 49, 29, 30, 31, 32, 46, 47, 33, 34, 35, 36, 38, 39, 40, 41, 42, 43]. Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g. [50, 51, 52, 53, 54, 55, 44, 56, 57]. Speaker characterization is the natural complement of speaker idenfication [13, 58]. Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming. Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research. With respect to privacy, it contains two obvious challenges for privacy. We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g. [59, 60, 61, 62, 63, 64, 65], or we can use natural language processing to anonymize the text content, e.g. [66, 67, 68, 69, 9, 70].
Privacy is even more important in always-on applications like wake-word detection, i.e. when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”. This application is more sensitive exactly because it is always on as users cannot choose when data is processed [71, 72, 73, 74]. The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [75, 76]. Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [72, 77, 78, 30, 79, 80, 81].
Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [82]. This task can have an impact on privacy in two primary ways. First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [83, 84]. Attenuating or removing background noises and competing speakers, as well as removing reverberation, can thus improve privacy. In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [80, 18, 85].
Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [86, 87, 13, 88, 89]. Some, however, limit the term biometric to refer only to physical and behavioral characteristics. Such properties include health [90], emotions [91] and gender identity [92, 93, 73, 46], each warranting their own treatment.
@@ II Content categories and applications
Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows). Each threat is marked with the number of the corresponding part of section IV-D. Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
III Definitions
III-A Privacy and security
@@
Privacy is closely related to challenges in security, and it is often difficult to distinguish between them. Here, information security refers to enforcing rightful access to information in scenarios where the access rights are binary (granted or not granted). In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [94]). The rules may distinguish between the ability to access and de facto accessing the data, different ways and purposes of using the data, and sharing the data with third parties. The rules may further be explicit (e.g., verbalized or documented) or implicit (as based on, e.g., social norms and tacit expectations) and can include requirements on monitoring, reporting, and control. For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig. 1). Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
III-B Information categories
@@
Target information is the content of a communication that is required to fulfill the purpose of said communication. Side-information (also known as auxiliary information, external knowledge, or background knowledge [95]) is any and all information in a communication other than the target information. See fig. 2.
@@
Attributes (or characteristics [4]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features. Table I lists examples of speech attributes with private information.
Observe that while the target information is often the lexical information contained in speech, in many applications, the target information can be some other attribute. For example, in medical analysis, the target information is the health state of the speaker. In that case, the lexical information may be private side-information.
@@
!! Target information Side-information
@@
Figure 2: The high-level threat model of speech interaction, where a target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message. It is a threat to privacy when an undesired recipient gains access to that target information or side information (marked by red arrows and exclamation marks).
III-C Tasks and processing objectives
When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer, and this is known as the trusted task. In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform other nefarious tasks. [47, 92].
Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with those of another. Anonymization of a speech signal refers to voice conversion where the purpose is to conceal private information. Pseudonymization further refers to reversible anonymization, where the original signal can be recovered if the concealed attributes are known. [^1] These definitions of anonymization and pseudonymization are outcomes of a community-wide discussion about personally identifying information (PII) at the Lorentz Center workshop “Speech as PII: Ethics & Usability in vis-à-vis with Security & Privacy” in 2021. The anonymized speech signal then corresponds to a speech signal of an artificial person, who is known as a pseudospeaker.
Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [96]. Re-identification refers to identifying individuals from apparently anonymized information [88, 97].
In anonymization tasks, enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s). It is used in the context of trusted tasks, where the agent is authorized to access this information. In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data. Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data. [98]
III-D Attack and attacker properties
We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences. A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [99, 100]. The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [101].
IV Threat and attack models
To better understand the decisive aspects of threats, it is useful to model threats and their components. The most important components of such models include the exposed data sources (section IV-A), tactics used by, and data available to the attacker (section IV-B), the flow of data on an abstract level, also known as the attack model (section IV-C), as well as the specific point in the system where private information leaves the intended path, known as the attack surface (section IV-D).
IV-A Vulnerable data sources
@@
Speech data can be exposed in two ways (see fig. 2): First, when a target information is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party. Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message. Typically, speech always contains consequential side information bundled into the target information. It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the target information. Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
@@
The difference between the two forms of information makes an important difference in approaches to mitigation. Transmitting the target information to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message. In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission for example through cryptography (see sections V-A, V-B and V-F). Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task. However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section V-A).
IV-B Attack types
@@
Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [102], where the attacker uses a known weakness on a large pool of potential victims. It is like spam email; if the attacker sends a large enough number of such emails, then randomly, someone triggers the trap by mistake. In this form of attack, anyone could be the victim. For example, a speech service could analyze private speaker attributes (for, e.g., health problems) or affiliations (with, e.g., minority groups), for purposes of targeted advertisement or surveillance. Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, birthday attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
@@
IV-C Attack model
In the design of privacy-protecting methods, attack models simulate and quantify an attacker’s performance. In other words, by putting themselves in the attacker’s place, safeguards designers can better understand pertinent aspects of weaknesses and quantify risks.
Figure 3 illustrates an attack model for measuring the extent of privacy [56]. We assume that there is some private data, which is anonymized such that it can be used in a trusted task. The attacker has access to the anonymized data, which is here called “public data”. Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it. The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information. Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
To complement an Attack(er) model, it can also be useful to define a Protector model for the use-case of the system. Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [103]. The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [104].
TABLE III: The Scenario-of-Use scheme [103].
Attacker Model (Attack on Privacy) Protector Model (Protection of Privacy)
Objective The information about the target speakers the attacker seeks to obtain. Objective Defense objective: The information in the spoken audio that must be protected.
Utility objective: The information in the spoken audio that must be retained.
Opportunity The attacker’s access to the target, including the availability of the target speakers’ audio. Also, access to (and any knowledge about) the protection. Opportunity The possibilities available for protection (i.e., counter-measures). Also, access to (and any knowledge about) the attack.
Additional
resources
The knowledge, data, and compute available to the attacker to carry out the attack (beyond the access to the target specified under Opportunity). Additional
resources
The knowledge, data, and compute available to the protector to defend against attack (beyond the access to the attack specified under Opportunity).
Attacks can be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used. Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker. For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics such as linguistic information. The anonymized sentences are known as trial utterances. The attack scenarios were then classified as [98]: